HIGH RESOLUTION FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE Ã^{2}Σ^{+} −X̃^{2}Π TRAN-SITION OF THE ICN^{+} ION

YUTAKA MIYAMOTO, ZDENEK ZELINGER, YOSHIHIRO NAKASHIMA and \underline{KEIICHI TANAKA}, Department of Chemistry, Faculty of Sciences, Kyushu University, Fukuoka, 812-8581 JAPAN.

The visible and near infrared emission spectrum in the 480 - 900 nm region of ICN^{+} was measured by Fourier transform spectrometer (Bruker IFS 120HR). The ICN^{+} was produced by Penning ionization of ICN with metastable He*. The ICN^{+} ion has much larger spin-orbit interaction constant (A = −4343cm^{−1}) than those of the ClCN^{+}(−276cm^{−1}) and the BrCN^{+}(−1477cm^{−1})^{a}. The Renner-Teller effect appears in the ν_{2} excited state of X̃^{2}Π. It is interesting to study the Renner-Teller effect of ICN^{+}. Eight vibronic bands of the Ã^{2}Σ^{+} − X̃^{2}Π transition (both for the Ω = 1/2 and 3/2 spin components) of ICN^{+} were observed. So far, the rotational analysis was performed for two vibronic bands, Ã^{2}Σ^{+}(000) − X̃^{2}Π_{3/2}(000) and Ã^{2}Σ^{+}(000) − μ^{2}Σ(010). Molecular constants, including the band origin, the effective rotational constant, centrifugal distortion constant and spin-rotation interaction constant, were determined for the Ã^{2}Σ^{+} and X̃^{2}Π_{3/2} states.

Due to the Renner-Teller effect, the X̃^{2}Π(010) state was split into four vibronic components, μ^{2}Σ, κ^{2}Σ, and ^{2}Δ_{P} (P = 3/2 and P = 5/2). For the Ã^{2}Σ^{+} − μ^{2}Σ band, P_{1}, P_{2}, R_{1} and R_{2}-branch were observed. For the μ^{2}Σ vibronic state, Ω-type doubling constant was determined as well as the rotational constant and centrifugal distortion constant. Renner parameter ϵ for the X̃^{2}Π state was determined to be −0.197 from the Ω-type doubling constant. The determined Renner parameter (ϵ = −0.197) was close to that of BrCN^{+}(−0.185). We are now analyzing the spectrum for the Ω = 1/2 spin component, Ã^{2}Σ^{+}(000) − X̃^{2}Π_{1/2}(000) and Ã^{2}Σ^{+}(000) − κ^{2}Σ(010), to study the Renner-Teller effect in more detail.

aJ. Fulara et al., J. Phys. Chem., 89, 4213-4219, (1985)